OpenStudy (anonymous):

Find the angle of elevation of the sun from the ground to the top of a tree when a tree that is 10 yards tall casts a shadow 14 yards long. Round to the nearest degree.

OpenStudy (anonymous):

Tan x=O/A Tan x=10/14 The angle B can be found using the inverse tangent function. Substitute in the values of the adjacent (14) and opposite (10) sides to angle B of the triangle. To convert radians to degrees, multiply by (180)/(`), since a full circle is 360^(o) or 2` radians. B=0.6202*(180)/(`) degrees Multiply 0.6202 by (180)/(`) to get (111.6449)/(`). B=(111.6449)/(`) degrees which is approximately equal to 3.1416. B=(111.6449)/(3.1416) degrees Divide 111.6449 by 3.1416 to get 35.5377. B=35.5377 degrees.

OpenStudy (anonymous):

Thanks i really need some help :)

OpenStudy (anonymous):

dats okay .. hope yu've understood :)